OpenStudy (phebe):

HELPSSS!! MEDALSS!!+FNSS!! What is the length of the hypotenuse of a right triangle if the two legs are 21 and 20?

OpenStudy (acxbox22):



OpenStudy (acxbox22):

a^2+b^2=c^2 20^2+21^2=c^2 400+441=c^2 841=c^2 c=sqrt841 c=29

OpenStudy (solomonzelman):

\[20^2+21^2=c^2\]

OpenStudy (phebe):

thanx

OpenStudy (acxbox22):

np

OpenStudy (phebe):

XD